NAOJ GW Elog Logbook 3.2


[Aritomi, Yuhang]
Here is information of lock of IRMC.
Input power: 2 mW, output power: 1.56mW (transmissivity: 78%)
Filter for lock of IRMC:
SR560 with low-pass 30 Hz, gain 10
40dB attenuation
We put an half-wave plate on the path before homodyne BS to get p-pol.
Then we measured reflection and transmission from BS for homodyne detector.
Reflection: 0.74mW, Transmission: 0.82mW
BS (BSW41-1064) somehow does not split the beam half and half now.

[Yuhang, Eleonora]
We have tried to recover the alignment and the lock of the filter cavity.
The alignment was easily recovered even if we could not use the local control for the telecope. See entry #1139. (We just tweeked a bit the BS position wiith picomotors)
We optimized the aligment with the input and output mirrors local control and it seems it is quite fine (flashes movie here:https://drive.google.com/open?id=1045raQj_n84CjnnUuYfYBT6Kqngv-id8) nevertheless we could not lock the cavity. We didn't have much time to investigate. Some possibilties:
1) The power is too large (a factor 3 higher than before), so the gain of the loop could be too high (we tried to change the gain value and also put some attenuation but it didn't work)
2) The beam fluctuation it too large since the BS and PR are not controlled (but this should be visible in the flashes).
3) Since the power is higher the camera in transmission could saturate more than before and mask the presence of HOM flashes. (We suspect there coud be a higher mismatching due to change of the beam from SHG)
The initial value of the gain were: input attenuaion = 5 and PZT gain = 4. As mentioned before we tried to change them but still we coudn't achieve the lock.
We will try to go back to the previouse power value and setting configuration to see if we can lock in this condition.

[Eleonora Yuhang]
While trying to recover the filter cavity alignement we found that the target PC used for the telescope local control had died.
We already had this issue which was solved by changing the power box. See entry #1035. We tried it again with another power box. It seemed to work fine and the supervisor PC could connect to it but we were not able to read and write any signal. We suspect a problem with the ADC/DAC unit connection. We will try another power box.

[Takahashi-san, Yuhang, Eleonora]
With the help of Takahashi-san, we have switched off all the instrumentation in TAMA in view of the power shutdown planned for tomorrow.
In particular we switched off the vacuum system in the central area and south arm and the compressor in the elecshop. We also closed the gate valves between the TM towers of the south arm and the pipe.
Some details of the swiching off precedure are reported in the attached file.

Participaint: Chienming, Shurong and Yuhang
Today Chienming and Shurong aligned the SHG, GRMC and OPO. There are some coments from them.
1. The alignment of SHG was always degraded by the constantly shifting of the second Faraday isolator.
2. There are two higher order modes of SHG cannot be removed. We guess these two modes come from some inherent mismatch inside SHG. We also found the bad shape of green beam comes from the cut of this green beam by the edge of SHG housing. However, as Chienming suggested, the side of beam is trivial compared with center. And we also found it is fine for filter cavity locking. So we guess it is fine.
3. Chienming put small spacers for the FI just after SHG. The purpose is to move the point(diffracting the light) to the edge of the beam.
4. GRMC and OPO were aligned to the best situation by Chienming and Shurong.
After that we found the beam was shifted. I guess this is because of the align of beam done by Chieming. Anyway, we recover the green beam direction by using BS before 100mm lens. Then we changed the AOM driving amplitude, the value now is -6.4dBm. It used to be -10dBm. This change of driving amplitude was suggested by Chienming. We increase driving amplitude and look at AOM first order power. We stop when we found the maximum. At that time, it became -6.4dBm. We can see from the attached figure. The zero order now is not round.
However, we still can see the beam seems to be cut by AOM. But Chienming suggested to check this beam with a much lower power(like what we have before, 8mW). But anyway, we cannot see this structure when there is only zero order going inside AOM.
So tomorrow we can do
1. Check the first order of AOM with lower power
2. Try to align filter cavity to see flash
Figure: (1and2: BAB inside OPO. 3and4: SHG. 5and6: GRMC 7:AOM situation of beam shape when amplitude maximized 8:AOM RF driving amplitude)

Participaint: Eleonora and Yuhang.
After the measurement of parametric ampkification, we considered to recover our filter cavity. But we found the beam was cut by EOM(GRMC/FC) and AOM. At the begining, we think we can recover the beam if we recover it for GRMC. However, it seems not like that.
Then we tried to align the beam better for EOM(GRMC/FC). We found the situation becomes better.
The entry 893 shows the spectrum of GRMC trasmission. I compared it with the GRMC situation now. I found there is a higher order mode is a little bit higher than before. Maybe this is the reason why we found the problem of beam cutting.
I am thinking maybe we can try to move the first lens(the lens just after SHG) to remove that peak.
Or tried to move 100mm lens just after the BS which seperates the beam going to GRMC and filter cavity.
Or we can further try to do the alignment of beam into AOM.

I summarize here some information that can be useful in order to design the telescope to match the squeezing beam to the filter cavity.
1) The distances from the last hole of the bench to the 2 inch telescope mirror (and the optics on this path) are reported in pic1. They have been measured by Yuefan in entry #441.
2) The target beam for the telescope is reported in the second attachment. The beam waist should be about 1 mm and it should be located close to the 2 inch telescope mirror (named M1 in the attached table and scheme). See also entry #442 for the former telescope design.

I aligned again the HeNe probe and made the calibration on the surface sample and on the bulk sample. Incident power= 35mW
I repeated the measurement on the small sapphire sample: a scan on the center, and a map at half thickness (according to the scan). Incident power 10W.
Then I made the ratio between my map and the one measured at SPTS as I did in entry 985. The ratio is now 1.17+/-0.2.

Participaint: Chienming, Shurong and Yuhang
Our measurement method: Injecting bright alignment beam(BAB) and pump beam inside OPO. Scanning phase with 500Hz but scanning OPO's PZT with only 5Hz. Then we use small amplitude of OPO's PZT to see the parametric gain effect inside the scanning peak. As you can see from the first picture.
Then by using this oscillated peak, we did the alignment of the pump beam into OPO. The improvement of alignment will bring more amplification for the peak of scanning. So we aligned the two steering mirrors between GRMC and OPO. After alignment, we get that peak as figure 2.
Then we tried to do this for pump power from 10mW to 79mW. For looking at parametric gain more clearly, we stop the scanning of OPO. On contrast, we manually tune OPO around resonance and use cursor to mark down the maximum and minimum of the oscillation. The result of these values is shown in tha attached figure 3.
Then we fit this result with the formula
gain = (1+B/Bth)^2/(1-B^2/Bth^2)^2 where B is the pumping field amplitude: P = B^2
We found a threshold of 80mW. We also clearly see the lasing when we provide 79mW of green pump.

[Shurong, Yuhang, Eleonora]
We have completed the assembly of the auxiliary IR mode cleaner, which will be used for the matching of the beams on the homodyne detector.
Fot the input/output mirros we used two mirrors from the same batch (6 pieces) used for the other IR modecleaner.
We double checked that all the mirrors were installed with the side arrow pointing towards the cavity
We put the white orings only on the input/output mirror, which should be fine.

Participaint: Aritomi, Chienming, Eleonora, Shurong and Yuhang
At the beginning, we want to do like this. Scanning green beam phase with a 1Hz, 1.5Vpp ramp signal. Then adjusting MZ offset to have different power going to OPO. Lock p-pol and adjust PLL to make both p and s-pol resonant. Locking OPO by p-pol. Sending green. Measuring OPO s-pol(BAB) transmission DC to see the parametric amplification and de-amplification.
1. Today's change: For making sure we are locking OPO at resonance. We set up a DC/RF photo detector for OPO s-pol transmission.
2. Lock OPO with p-pol, found demodulated s-pol transmission RF oscillating at 1Hz. The RF signal before demodulation oscillates at 15MHz with an amplitude modulated at 1Hz.
3. The green injection changes the co-resonant condition. We guess this is because of temperature change. See attached picture 1.
We did like this finally. Manually put the transmission of the seed beam on maximum pk-pk. Don't use p-pol lock and PLL. For different pump power, we measured the amplification and deamplification. See attached figure 2. We did the fit and found the threshold now should be around 140mW.
Since we know the green changed the temperature of OPO crystal, we also checked the temperature influence. For a green power of 60mW, we changed temperature from 6.9kOm to 7.2kOm. Now the optimal temperature seems to be 7.06kOm, which is 0.04kOm far from previous optimal temperature 7.02kOm. However, the influence of temperature change caused by green power seems not crucial. In any case, the non-linear crystal is still working in the broad optimal region.
However, we decided to check the matching.


Since we could not find any spare half-wave plate for 532 nm, we have temporarily removed that placed just after the last green faraday isolator, before the filter cavity.
We use it to operated the green mode cleaner in p-pol. The value of the angle in the original configuration is shown in pic 2.
We will need to buy (or find) some new ones.

We did more detailed test to the board last week.
Since last week when we powered up the board with NIM rack, it heated up very soon and had some bad smell. Just in case the NIM rack we have doesn't give the right power and may cause some damage, this time we powered up the board with external power supply. The first problem we found out, it is the 8V power supply of this board which give power to drive the galvo, has a much large current than it is supposed to be, that is also the reason why they board heat up so soon we guess.
At the beginning, we just powered up the 20V part of the circuit, and sent a sine wave instead of the quadrant signal to A,B,C,D port one by one. By checking each check-point inside the board, we could see clear sine wave.
Then we powered up also the 8V part, the signal at the check-point started to have high frequency oscillation in MHz range. We traced it back, found out the oscillation was already there since the beginning of this path.
When we wanted to do more investigation of the problem, we were looking for some checking point on the board according to the design, but we could not find them. So we checked the date written on the board, it has been done in Aug 1999, but the design was finshed later than that, so we guess there may be newer version of this board in TAMA and they could have better performance than the one we have in Nikhef now.

[Yuhang, Eleonora]
We found four more boards for the galvo control. Two of them were in the shelves of TAMA entrance, the other two (those on the right in pic1) were installed in one of the TAMA injection rack
All the board are dated Aug 1999 (see pic 3) but one reports the writing "corrected by Koji Arai & Sa 2002/12/10 " (see pic 2)
See entry #831 for more details about galvo and boards.

Participaint: Matteo and Yuhang
By considering the space and alignment of homodyne detector(HMD). We designed the telescope for both beam coming from local oscillator and bright alignment beam(BAB).
The attached figure 1 and 2 show the telescope. The attached figure 3 shows the overall situation in the future implementation.

Participaint: Aritomi, Chienming, Eleonora, Shurong and Yuhang
Today we achieved the lock of OPO on both p and s pol. The method is to lock OPO with p-pol (PDH signal goes to PZT of OPO) and then use PLL to lock s-pol to p-pol. Actually, before locking PLL we made sure the p and s both resonant inside OPO with peak overlaping.
The tricky part is to lock PLL as following:
1. We found that the frequency diffrerence of p and s is ~360MHz while the main laser current is 1.4A. However, this frequency increased ~420MHz after we increased main laser current to 1.8A.
2. Locking PLL for 420MHz is not possible. According to datasheet, we can provide VCO frequency only below 400MHz. At this moment, we decided to not use optimal temperature(optimal is 7.2kOm) but use 7.171kOm.
3. We investigated quite a long time and found we can use the built in divider to lock beatnote smaller than 400MHz. Since the phase detector can hold frequency only below 104MHz(also from datasheet), so we can divide signal and make it smaller than 104MHz. Then we can use reference frequency to compare with it.
4. As we have already mentioned in the last entry, we can only lock one side of PLL. We further found that we can change the sign of correction signal inside PLL board. We need to find the good comnbination of correction signal sign and the original frequency side(lower or higher than locking frequency). However, we were think if there is a way to change the offset of correction signal and make it have both positive and negitive signal.
After locking PLL, it is easy to lock OPO. The locking parameters are single-pol low-pass at 300Hz, gain of 100 and invert for SR560.
The locking parameters are single-pol low-pass at 30Hz, gain of 2 for SR560.(for GRMC)
I did the simulation about how much green we can produce now. The temperature we lock is 7.171kOm. The green we produce should be 90% of the peak value. However, the green production will be even smaller if we consider the overlap of p and s is not perfect.

The design of the QPD telescope has been finished.
The idea is to seperate the reflected beam from the cavity into two path, and put one QPD on each path, in a way that the guoy phase on the two quadrants has around 90 degree difference. So one of the quadrant could sense the shift and one could sense the tilt.
The target beam is decided by the size of the quadrant sensor :
- The size of each quadrant element is 11 mm2
- The gap between the elements is 70um
So I chose the target beam size at QPD should be around 1~1.5mm
The lenses I used in the simulation are the green lenses I could found on Thorlabs website.
In both of the case, the origin is set at the beam extracting port of the FI.
1.
Convex lens(mm) |
Concave lens(mm) |
Distance between origin and the first lens(cm) |
Distance between two lenses(cm) |
Beam size at QPD(um) |
Distance between concave lens and QPD(cm) |
500 |
-75 |
17.5 |
43 |
1300 |
105 |
2.
Convex lens(mm) |
Concave lens(mm) |
Distance between origin and the first lens(cm) |
Distance between two lenses(cm) |
Beam size at QPD(um) |
Distance between concave lens and QPD(cm) |
400 |
-50 |
42 |
42 |
1100 |
30 |
The guoy phase difference at two QPD is around 98 degree. So I think we could go on with this design.

Participaint: Aritomi, Eleonora, Chienming, Shurong and Yuhang
1. We checked the spectrum analyzers by signal generator. We found out
HP E4411B (the new one): Has a wrong amplitude value but higher bandwidth up to 1.5GHz.
HP 8563E(the old one which was used by Marco): Has a correct amplitude but the bandwidth is limited to 170MHz.
So we decide to check firstly by the new one and then tune beatnote frequency smaller than 170MHz. FIncally check amplitude by the old one.
2. The lesson we got from PLL locking
We need to load Marco set-up and remember to write this set-up into chip. How to change this set-up and implement it for the second board still needs to be investigated.
We checked the correction of PLL, we found now it is only positive. So we can bring initial frequency locked only when it is lower than reference frequency.
To do list:
1. investigate the amplification factor of each channel.
2. investigate the locking peformance of PLL(the highest locking frequency).
3. lock PLL to make both p-pol and s-pol resonant.
4. investigate how to apply lock for PLL board two.